I can't solve this problem, can you help me? Your correct answer will help me understand what I am doing wrong. Thank you.
-8n-8(-4-2n)

Answer by oberobic(2304)

Well, it is not an equation unless there is an '=' sign.
.
The expression can be simplified, but it cannot be solved.
.
-8n -8(-4-2n) = -8n +32 +16n = 8n + 32 = 8(n+4)
.
The point of a problem like this is perhaps to reinforce the concept that a negative times a negative is a positive.
.
Done.